The length of a rectangle is 13 cm more than the width. The perimeter is 65. Find the dimensions.
Answer by macston(5194)

P=perimeter=65cm; W=width; L=length=W+13cm
P=2(L+W) Substitute for L
P=2(W+13cm+W)
65cm=2(2W+13cm)
65cm=4W+26cm Subtract 26cm from each side.
39cm=4W Divide each side by 4.
9.75cm=W ANSWER 1: The width is 9.75cm.
L=W+13cm=9.75cm+13cm=22.75cm ANSWER 2: The length is 22.75 cm.
CHECK:
P=2(L+W)
65cm=2(22.75cm+9.75cm)
65cm=2(32.5cm)
65cm=65cm